Suppose X and Y are 2 independent normal random variable such that x~N(65,28) and Y~N(85,36).Find p(134<X+Y<166) and p(Y>x)


1
Expert's answer
2021-08-19T15:20:32-0400

The distribution of a sum of two normally distributed independent variables "X\sim N(\mu_1, \sigma_1^2)" and "Y\sim N(\mu_2, \sigma_2^2)" is another normal distribution "(X+Y)\sim N(\mu_1+\mu_2, \sigma_1^2+\sigma_2^2)."

Given "X\sim N(65, 28), Y\sim N(85, 36)." Then "(X+Y)\sim N(65+85, 28+36)."


"P(134<X+Y<166)"

"=P(X+Y<166)-P(X+Y\leq134)"

"=P(Z<\dfrac{166-150}{\sqrt{64}})-P(Z\leq\dfrac{134-150}{\sqrt{64}})"

"=P(Z<2)-P(Z\leq-2)"

"\approx0.97725-0.02275=0.9545"

The distribution of a difference of two normally distributed independent variables "Y\sim N(\mu_2, \sigma_2^2)" and "X\sim N(\mu_1, \sigma_1^2)" is another normal distribution "(Y-X)\sim N(\mu_2-\mu_1, \sigma_1^2+\sigma_2^2)."

Given "X\sim N(65, 28), Y\sim N(85, 36)." Then "(Y-X)\sim N(85-65, 28+36)."


"P(Y>X)=P(Y-X>0)=1-P(Y-X\leq0)"




"=1-P(Z\leq\dfrac{0-20}{\sqrt{64}})=1-P(Z\leq -2.5)"

"\approx0.9938"